Laverne buys a call option on XOM with a strike price of $90.00. XOM is already trading at $91.00 per share. Laverne pays a $1.65 premium on the call, which has a three month expiration. If XOM stock goes up to $103.00 per share and Laverne sells the call at a premium of $12.65, how much total profit will Laverne make on this transaction?

$126.50

$12.65

$1,265.00

$1,100.00

Shirley believes that Nike (NKE) stock is going to decline in value because of softening demand for Nike apparel (Under Armor is taking all of Nike's business). Shirley decides to buy a put on NKE with a strike price of $90.00 and an expiration of six months. NKE currently trades at $91.00. The premium on the put is $1.60. NKE increases to $101.50 per share at the end of the six month period. Shirley is still holding the put option at the end of the expiration period. Calculate Shirley's total gains or losses on this transaction at the expiration date.

Gain of $10.50.

Gain of $1,050.00.

Loss of $1.60.

Loss of $160.00.

Lenny decides that he wants some extra income, so he writes a call on McDonald's (MCD) stock with a strike price of $95.00. MCD is currently trading at $92.00. Lenny collects a premium of $0.50 on the transaction. If MCD stock goes up to $110.00 what will be the result for Lenny?

Lenny gains $15.00.

Lenny gains $14.50.

Lenny gains $1,450.00.

Lenny loses $1,450.00.

Squiggy writes a put option on Home Depot (HD) stock. The stock is currently trading at $108.00 per share. Squiggy writes a put option with a strike price of $105.00 and an expiration of 6 months. Squiggy collects a $1.05 premium. Assume the stock price collapses to $31.00 per share by the end of the expiration period (and results in a premium on the put option of $74.00). Calculate Squiggy's total gains or losses on this transaction.

Squiggy gains $74.00.

Squiggy gains $740.00.

Squiggy gains $7,400.00.

Squiggy gains $7,295.00.

Squiggy loses $7,295.00.
Answer #1

1.
=(Selling price-Purchase price)*contract size
=(12.65-1.65)*100=1100

2.
Profit=(MAX(strike price-spot price at expiry,0)-premium)*contract size
=(MAX(90-101.50,0)-1.60)*100
=-160
Loss of 160

3.
Profit=-(MAX(spot price at expiry-strike price,0)-premium)*contract size
=-(MAX(110-95,0)-0.50)*
=-1450
Lenny loses $1450

4.
Profit=(Purchase price-Selling Price)*contract size
=(1.05-74)*100
=-7295
Squiggy loses $7295